OpenStudy (anonymous):

the vertex of this parabola is at (2,-4). When the x-value is 3, the y-value is -1. what is the coefficient of the squared expression in the parabolas equation

OpenStudy (freckles):

so write like this y=ax^2+bx+c first

OpenStudy (freckles):

the coefficient is then what is in front of x^2

OpenStudy (freckles):

so the coefficient of x^2 for this particular parabola's equation I have written is a

OpenStudy (freckles):

so example: \[\text{ find coefficient of } x^2 \\ y=(3x+2)^2 \\ y=(3x)^2+2(2)(3x)+(2)^2 \\ y=9x^2+12x+4\] The coefficient of x^2 is 9

OpenStudy (freckles):

And if you wanted to find the coefficent of x you would say 12 in this case

OpenStudy (freckles):

and if you wanted to know the constant term it is 4 in this case

OpenStudy (freckles):

in general, once you have in this form y=ax^2+bx+c then you can determine the following: coefficient of x^2 is a coefficient of x is b constant term is c

OpenStudy (freckles):

oh I didn't see that whole question

OpenStudy (freckles):

I just seen how to find the coefficient of a x^2 for a parabola sorry

OpenStudy (freckles):

\[y=a(x-h)^2+k \text{ where } (h,k) \text{ is the vertex}\]

OpenStudy (freckles):

what is h? what is k?

OpenStudy (freckles):

hint: h and k are given in your problem

OpenStudy (freckles):

well to find a you will then use f(3)=-1 <--but do this part after pluggin in h and k

OpenStudy (freckles):

a will be your answer

OpenStudy (freckles):

when you find it

OpenStudy (freckles):

let me know if you come back and need my help